OpenStudy (anonymous):

Solve the system by substitution. {2.5x - 3y = -13 {3.25x - y = -14 I usually have no problem with the substitution method, but this one is stumping me.

OpenStudy (anonymous):

I started working on it, then got stuck.

OpenStudy (anonymous):

solve the bottom equation for y y = 3.25x+14 plug y into the top equation. solve for x. then plug the x value into either equation to get y.

OpenStudy (anonymous):

I started it that way by solving for y. I multiplied the y equation by -3 to cancel out the y

OpenStudy (anonymous):

but then somwhere after that it got messed up.

OpenStudy (smokeysthename):

are you still stuck on it?

OpenStudy (anonymous):

yes

OpenStudy (smokeysthename):

ok gimme a sec and lemme work it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3.25x -y = -14 y = 3.25x+14 2.5x-3y = -13 2.5x-3(3.25x+14)= -13 2.5x-9.75x-42= -13 -7.25x = 29 x= -4

OpenStudy (smokeysthename):

tomo's explaining it.

OpenStudy (smokeysthename):

plug x back into the original equation of: {3.25x - y = -14

OpenStudy (anonymous):

oh ok, I did it wrong then I think. I started with 3.25x - y = -14 and multiplied the equation by -3; -3(3.25x - y = -14) = -9.75x +3y = 42

OpenStudy (smokeysthename):

yeah thats solving by elimination is what you're trying to do.

OpenStudy (anonymous):

ohhhh oops. so x = -1

OpenStudy (smokeysthename):

tomo said that x = -4

OpenStudy (anonymous):

I I mean y = 1 and x = -4

OpenStudy (anonymous):

y = 1 and x = -4

OpenStudy (smokeysthename):

yes thats correct (:

OpenStudy (smokeysthename):

good job!

OpenStudy (anonymous):

yay thanks!

OpenStudy (smokeysthename):

no problem! :D